# UniMD: Towards Unifying Moment Retrieval and Temporal Action Detection

### v1

@@ Abstract
Temporal Action Detection (TAD) focuses on detecting pre-defined actions, while Moment Retrieval (MR) aims to identify the events described by open-ended natural language within untrimmed videos. Despite that they focus on different events, we observe they have a significant connection. For instance, most descriptions in MR involve multiple actions from TAD. In this paper, we aim to investigate the potential synergy between TAD and MR. Firstly, we propose a unified architecture, termed Unified Moment Detection (UniMD), for both TAD and MR. It transforms the inputs of the two tasks, namely actions for TAD or events for MR, into a common embedding space, and utilizes two novel query-dependent decoders to generate a uniform output of classification score and temporal segments. Secondly, we explore the efficacy of two task fusion learning approaches, pre-training and co-training, in order to enhance the mutual benefits between TAD and MR. Extensive experiments demonstrate that the proposed task fusion learning scheme enables the two tasks to help each other and outperform the separately trained counterparts. Impressively, UniMD achieves state-of-the-art results on three paired datasets Ego4D, Charades-STA, and ActivityNet. Our code will be released at https://github.com/yingsen1/UniMD.
@@ 1 Introduction
Temporal Action Detection (TAD) and Moment Retrieval (MR) are two similar tasks that seek to identify specific events and their corresponding temporal segments within untrimmed videos. TAD focuses on identifying temporal segments related to a single action, while MR aims to identify temporal segments that align with a natural language description. These tasks have significant connections, for instance, most natural language descriptions in MR are a combination of multiple actions from TAD. However, most existing works [64, 8, 26, 66, 58] treat TAD and MR as two independent tasks, employing separate models for each. In this paper, we aim to answer a question: whether the two tasks can benefit each other by fusing them using a single model?
Figure 1: Illustration of TAD, MR, and UniMD. UniMD performs both TAD and MR in a single model (i.e. same model parameters).
Figure 2: Co-train results evaluated on ActivityNet under various data volumes in the training set. Dedicated models are trained separately for specific tasks, using overall videos as training data. When co-trained using a fraction of the training data, it can even achieve superior performance compared to dedicated models, such as 25% training data in MR and 50% in TAD.
@@
To investigate such potential synergies between TAD and MR, we propose a new task formulation termed Moment Detection (MD) that aims to address both TAD and MR tasks simultaneously, as illustrated in Fig. 1. For MD, we design a task-unified architecture, Unified network for Moment Detection (UniMD), with a task fusion learning approach, to enhance the performance of both tasks:
@@
(2) Task fusion learning. To further improve the synergies between TAD and MR, we explore task fusion learning to promote the mutual influence of both tasks. Specifically, we examine and discuss the influence of pre-training and co-training in task fusion learning. Moreover, we introduce two co-training methods, namely synchronized task sampling and alternating task sampling. Synchronized task sampling prioritizes video samples that encompass both tasks, ensuring that each training iteration includes both tasks. In contrast, alternating task sampling updates the network based on an alternating task at each iteration. Among them, the co-training with synchronized task sampling effectively enhances the synergy between the two tasks and leads to distinct improvements for each task. As illustrated in Fig. 2, the proposed co-trained model can achieve better results compared to dedicated models, even with only a subset of training data, i.e. , 25% training videos for the MR task and 50% training videos for the TAD task. This demonstrates that the mutual benefits are not merely derived from the increased quantity of annotations, but rather from the enhanced effectiveness of co-training.
To summarize, this paper makes three contributions: (i) we propose a unified framework, termed UniMD, to address the tasks of TAD and MR simultaneously. It establishes the relationship between actions and events by inheriting the text encoder of CLIP to encode the query and predicts the action- or event-relevant temporal boundaries through the query-dependent heads. (ii) To the best of our knowledge, we are the first to exploit task fusion learning to investigate whether the two tasks can help each other, and propose an effective co-training method to enhance their synergy, which provides valuable insights and experiences for the community on training large language-based video model (e.g. LLaMA-vid [28]) on action/event detection in the future. (iii) Extensive experiments demonstrate that UniMD achieves the state-of-the-art performance across various benchmarks, namely, 23.25% mAP in Ego4D-MQ, 14.16% R1@30 in Ego4D-NLQ, 63.98% R1@50 in Charades-STA and 60.29% mAP@50 in ActivityNet.
@@ 2 Related Work
Temporal Action Detection (TAD). TAD is dedicated to the identification of predefined actions and their temporal segments in untrimmed videos. It can be classified into two-stage methods [62, 71, 69] and one-stage methods [64, 29, 31, 44]. Two-stage methods initially generate video segments as action proposals, which are subsequently categorized into actions and refined in terms of their temporal boundaries. Conversely, one-stage methods aim to localize actions within a single shot without relying on action proposals. Some TAD tasks involve a substantial number of action categories densely labeled in videos. To tackle this challenge, several approaches [8, 43, 9, 10] have been developed to strengthen the dependencies between actions and capture both short-term and long-term dependencies.
Moment Retrieval (MR). In contrast to TAD, MR aims to identify specific actions using open-ended natural language descriptions. MR can be categorized into one-stage and two-stage methods. Two-stage methods [15, 65, 60, 67], also known as proposal-based methods, utilize sliding windows and proposal generation models to generate a set of candidate windows. These candidate windows are then ranked, and the one that exhibits the highest similarity to the given description is selected as the final result. One-stage methods [61, 63, 38, 34], also referred to as the proposal-free method, directly compute the time pair or output confidence scores for each video snippet. Since MR searches for open-ended events rather than closed-set categories, it is crucial to fully utilize the semantic information within the queries. Therefore, several methods [61, 38, 7, 66] focus on incorporating multi-modal features and enhancing cross-modal interactions to improve boundary regression.
Vision-Language Modals. Studies on the interaction of vision and languages have been ongoing for several years. Recently, some methods [42, 19, 59] scale up training processes by leveraging large-scale datasets, resulting in the progress of robust visual-text alignment. Furthermore, many efforts [52, 55] have been made to extend these methods to video-text modals. Consequently, an increasing number of researches [72, 57, 24, 56, 6, 27, 17] focus on effective techniques for fusing vision-language representations and utilizing them to establish connections across various tasks, aiming for a universal objective.
@@ 3 Problem Definition
TAD focuses on identifying temporal segments of a pre-defined set of action categories in untrimmed videos, as shown in the first row of Fig. 1. Formally, given $C$ action categories, a TAD function $f_{\text{tad}}$ takes video images/features $X$ as input and generates the temporal segments of the action categories:
@@
where $s$ and $e$ denote the start time and end time of the action category $c_{\text{tad}}\in[1,C]$. Different from TAD, as shown in the second row of Fig. 1, MR aims to predict the temporal segments of the events described by the open-ended natural language:
@@ 4 Task-unified Architecture
where $Q=\left\{q_{i}|i\in[1,C+N]\right\}$, and $conf\in\mathbb{R}^{C+N}$ denotes the foreground confidence corresponding to the close-set action categories from TAD or the events from MR.
@@ 4.1 Encoder
where $Z=\left\{Z^{1},Z^{2},\dots,Z^{L}\right\}$ and $L$ denotes the number of feature pyramid levels.
@@ 4.2 Decoder
$\left\{conf^{t}\mid t\in[1,T]\right\}&=\Phi_{\text{cls}}\left(G,Z\right)\\ &=\Phi_{\text{sigmoid}}(\Phi_{\text{scale}}(G\odot\Phi_{\text{conv}}(Z))))),\\ \end{aligned}$
@@
$\left\{(d_{\text{s}}^{t},d_{\text{e}}^{t})\mid t\in[1,T])\right\}&=\Phi_{\text{reg}}(G,Z)\\ &=\Phi_{\text{relu}}(\Phi_{\text{scale}}(\Phi_{\text{q-conv}}(\Phi_{\text{conv}}(Z)))),\end{aligned}$
@@ 6.2 Implementation Details
Network. We use the text encoder from CLIP-B [42], with frozen parameters during training. The vision encoder includes seven ConvNext blocks. Within them, two blocks serve as a stem, while the remaining five layers 2$\times$ downsample to generate the feature pyramid. The FPN incorporates three layers of BiFPN [48]. In decoder, the classification head comprises two convolutional layers and one query-dependent convolutional layer. The regression head follows the same configuration, and includes a query transformation branch that incorporates three fully connected (fc) layers with a dimension of 512. In both classification head and regression head, the dimension of the first two convolutional layers is set to 512. Other parameters, such as the regression range for feature pyramid, center sampling strategy, and EMA model [18], are set according to [64].
@@
Training. The training of three paired datasets utilizes a consistent AdamW optimizer and follows the same loss weights ($\lambda_{tad}=3$, $\lambda_{mr}=1$). However, there are slight differences in other settings. For Ego4D, the model is trained for 15 epochs with a warm-up period of 5 epochs. The batch size is set to 2 and the learning rate is set to 1e-4. In the case of Charades, the model undergoes 20 training epochs, with 10 epochs as a warm-up, a batch size of 2, and a learning rate of 1e-4. For ANet, it is trained for 15 epochs with an initial learning rate of 1e-3, a batch size of 8, and a warm-up period of 5 epochs.
@@ 6.3 Ablation Study
[…]
In this section, we conduct ablation experiments on the proposed query-dependent head, the task fusion learning, loss weights of two tasks, and data volumes in the training set.
Query dependent regression head. As shown in Table 1, the regression head with queries as input yields better performance for both tasks (21.86$\rightarrow$22.61mAP in TAD and 12.49$\rightarrow$13.99R1@30 in MR). This enhancement is attributed to the semantic information of the query, which enables more precise time intervals for specific actions. In the query transformation branch of this head, the MLP that is equipped with 3 fc layers achieves the best performance.
@@
Table 2: Effective of task fusion learning evaluated on the Ego4D validation set. For simplicity, synchronized task sampling and alternating task sampling are abbreviated as “Sync.” and “Alt.”.
[…]
@@
Table 3: Effect of loss weights on validation of Ego4D, using synchronized task sampling.
Loss weights. The impact of different loss weights on the outcomes is significant, as indicated by Table 3. For example, when using a “1-1” loss weight, the co-training performance is inferior to the dedicated model for TAD (22.55 vs. 22.61 mAP). Therefore, we carefully choose a loss weight of “3-1” to ensure that the co-train models are trained with a suitable balance of tasks.
@@
Table 4: Impact of various ratios of data volumes evaluated on validation of ANet and val_2 of ANet-Caption.
Data volumes. The impact of data volumes on the validation of ANet is presented in Table 4, By controlling the number of training data for each task (e.g. 0.5-0.5 means only half of the videos from TAD and MR), we observe a positive effect on both tasks with an increase in data volumes, as shown in Fig. 2. Noteworthily, in TAD, the co-trained model using only 50% of the training videos slightly outperforms the dedicated model, while in MR, even using only 25% of the training videos significantly surpasses the dedicated model. It reveals that the mutual benefits are not merely derived from the increased quantity of annotations, but rather from the enhanced effectiveness of co-training.
@@ 6.4 Comparison with the state-of-the-art
Validation
EgoVLP (VSGN)[30] E 10.69 29.59 - - - -
InternVideo (AF)[5] Vverb 20.69 37.24 - - - -
@@
EgoVLP (VSLNet)[30] E - - 10.84 6.81 18.84 13.45
InternVideo (VSLNet)[5] Vverb - - 12.32 7.43 21.73 15.07
@@
UniMD+Sync. Vverb 24.12 44.14 14.48 10.09 35.00 25.04
Test
EgoVLP (VSGN)[30] E 9.78 27.98 - - - -
InternVideo (AF)[5] Vverb 19.31 35.58 - - - -
@@
UniMD+Sync. Vverb 23.25 44.80 14.16 10.06 26.95 19.16
Table 5: Comparison with the state-of-the-art in TAD and MR tasks on Ego4D validation set and test set. In Ego4D, Moment Query is TAD task, and Natural Language Query is MR task. The feature E represents EgoVLP, the Vverb denotes VideoMAE-verb, the Vnoun represents VideoMAE-noun and the Fens refers to ensemble features.
@@
MomentDETR[26] SF+CLIP - 53.63 31.37 - -
MomentDETR(pt)[26] SF+CLIP - 55.65 34.17 - -
QD-DETR[37] SF+CLIP - 57.31 32.55 - -
@@
Table 6: Comparison with state-of-the-art methods evaluated on test set of Charades for TAD and Charades-STA for MR. The SF denotes SlowFast.
@@
Table 7: Comparison with state-of-the-art on the validation set of the ANet for TAD and val_2 set of ANet-Caption for MR. The SF denotes SlowFast. The features used in ${}^{{\dagger}}$ cannot be obtained.
@@ 7 Conclusion
This paper provides an answer to the question of “whether the TAD and MR tasks can benefit each other by fusing them in a single model?”. To achieve this, we begin by designing a task-unified network, termed UniMD, with uniform interfaces for task input and output. We further explore various task fusion learning methods to enhance the collaboration between TAD and MR. Through experiments on three paired datasets, we present evidence that the task fusion learning approach effectively improves the performance of both tasks.
Supplementary Material
Temporal Action Detection Moment Retrieval
#clip #$\text{\#clip}_{\text{diff}}#$cls $\text{Avg}_{\text{dura}}$(s) $\text{Avg}_{\text{ins}}$ $\text{Avg}_{\text{ins\_dura}}$(s) #$\text{\#clip}_{\text{overlap}}#$clip #$\text{\#clip}_{\text{diff}}#$query $\text{Avg}_{\text{dura}}$(s) $\text{Avg}_{\text{ins}}$ $\text{Avg}_{\text{ins\_dura}}$(s)
Ego4D
train 1.5k 1.5k 17 1k 1k 11.3k
val 0.5k 0.5k 7 0.3k 0.3k 3.9k
total 2k 2k 110 472 9.2 42.8 24 1.3k 1.3k 15.2k 494 10.2 10.6
Charades
train 8k 2.6k 5.3k 5.3k 0 12.4k
val 1.8k 0.5k 1.3k 1.3k 0 3.7k
total 9.8k 3.1k 157 30 6.8 12.9 6.6k 6.6k 0 16.1k 30 2.4 8.3
ActivityNet
train 1w 15 9972 1w 37 3.7w
[…]
val_2 41 4885 0.5w 0 1.7w
total 2w 65 200 118 1.5 49 19774 2w 37 7.1w 118 3.5 35.5
Table A: Statistics of three paired datasets that cover both TAD and MR tasks. #clip denotes the number of video clips, #$\text{\#clip}_{\text{diff}}#$ presents the number of non-overlapping videos, and #$\text{\#clip}_{\text{overlap}}#$ refers to the number of overlapping videos. $\text{Avg}_{\text{dura}}$, $\text{Avg}_{\text{ins\_dura}}$, and $\text{Avg}_{\text{ins}}$ indicate the average video duration, the average instance duration, and the average number of instances per video, respectively. #query represents the number of queries in MR while #cls denotes the number of action categories in TAD. The number underlined in ActivityNet indicates that the validation set in TAD only has one version, while ActivityNet-Caption has two versions, val_1 and val_2.
8 Appendix
In the appendix, we describe (1) more details about three paired datasets (Section A), (2) additional ablation experiments (Section B), and (3) qualitative results (Section C). For sections, figures, tables, and equations, we use numbers (e.g. Sec. 1) to refer to the main paper and capital letters (e.g. Sec. A) to refer to this appendix.
Appendix A Details About Datasets
Statistical analysis of datasets. (1) Ego4D [16] is a large-scale first-person video dataset that consists of $\sim$3000 hours of egocentric videos. The TAD task of Ego4D includes 110 action categories. Specifically, the TAD task has an average of 9.2 instances per video across 2k videos, while the MR task has an average of 10.2 instances per video across 1.3k videos. (2) Charades and Charades-STA. Charades [46] is a large-scale dataset that consists of 9.8k videos, with an average duration of 30 seconds. The dataset includes a total of 157 action categories that are densely labeled in the videos. Charades-STA [15], an extension of Charades, introduces natural language descriptions as annotations. Charades-STA consists of 6.6k videos, with the same average duration as Charades. (3) ANet and ANet-Caption. ActivityNet (ANet) [3] contains 200 action categories and $\sim$20,000 videos. Additionally, ActivityNet-Caption (ANet-Caption) [25] is built upon ANet, including descriptions of events and their corresponding timestamps. ANet-Caption also comprises $\sim$20,000 videos, having the same average duration of 118 seconds as ANet.
However, the proportion of overlapping videos between the TAD and MR tasks varies across datasets, as shown in Table A. In Ego4D, there is a minimal overlap of only $\sim$2%. Conversely, ANet exhibits a substantial overlap, with more than 99% of the videos being shared. In the Charades, all of the videos from Charades-STA are included, while Charades-STA covers 67% of the videos in Charades. Since each pair of datasets shares the same video source, the duration of the videos in TAD is approximately identical to that in MR. In addition, there are notable differences in the number of instances and average instance duration. In Ego4D, despite having similar average instance numbers, there is approximately a 300% difference in average instance duration. In Charades, TAD has an average instance number 2.8 times that of MR, and its average instance duration is 50% longer than that of MR. ANet presents a different situation, where MR has a higher average instance number compared to TAD, but the average instance duration is shorter in MR.
Annotation visualizations. Figure A displays six additional annotation visualizations that illustrate the relationship between TAD and MR.
Appendix B Additional Ablation Experiments
In this section, we explore the model settings of the vision encoder and decoder based on TAD metrics of Ego4D, as presented in Table B. Additionally, we investigate three solutions for ANet (Table C) and analyse the effect of task fusion learning on Charades (Table D) and ANet (Table E).
Component TAD
Encoder #$\text{\#}D_{reg}$ ${scale}_{cls}$  mAP  R1@50
128 21.86 38.79
256 22.19 40.62
512 Y 22.61 41.18
ConvNext 512 N 22.05 38.58
Transformer 512 Y 21.40 38.90
Table B: Effect of different network components evaluated on the TAD task of Ego4D validation set under individual training. The TAD task is the Moment Query task in Ego4D. The #$\text{\#}D_{reg}$ means the dimensions in the query transformation branch of the regression head and the ${scale}_{cls}$ indicates the scale operation in the classification head.
The choice of encoder and decoder. For the ablation study on the vision encoder, we conduct an experiment by replacing the ConvNext [35] blocks with Transformer units, following the same pipeline. As shown in Table B, using ConvNext as the main block of the vision encoder yields better performance (22.61% vs 21.40% in mAP) compared to Transformer. Additionally, we conduct an exploration of the query transformation branch in the regression head and the learnable scaling operation in the classification head. The comparison indicates that using a Multi-Layer Perceptron (MLP) with a channel size of 512 leads to superior semantic translation, resulting in a higher mAP. Furthermore, the integration of learnable scaling operations positively affects performance.
Method solution TAD MR
 mAP  mAP@50  R5@50  R5@70
MR - - - 77.28 53.86
TAD 1 37.78 57.48 - -
TAD+MR 37.98 57.48 79.90 53.71
TAD ${}^{{\dagger}}$ 38.60 58.31 - -
TAD+MR 39.82 60.04 80.83 57.21
TAD ${}^{{\dagger}}$ 38.16 57.42 - -
[…]
Table C: Effect of the three solutions for TAD evaluated on the validation set in ANet and val_2 split in ANet-Caption. Here, the methods referred to as “TAD+MR” utilize task fusion learning with random task sampling. The methods “TAD” and “MR” are dedicated models for each task. Solutions ${}^{{\dagger}}$ and ${}^{{\dagger}}$ utilize score fusion strategy.
Three solutions for ANet. In the context of score fusion strategy in ANet, we explore three solutions for TAD, as outlined in Sec. 6.2. Based on the comparison presented in Table C, the score fusion strategy, implemented in solutions 2 and 3, enhances the performance of TAD compared to solution 1 which does not use external classification scores. Besides, among the results of task fusion learning, this approach offers an advantage. Specifically, solution 2 which averages the text embeddings of entire categories, shows slightly better performance (mAP and mAP@50 in TAD, and R5@50 in MR) compared to the manual prompt for action proposals in solution 3.
Method policy TAD MR
 mAP  R1@50  R1@70  R5@50  R5@70
TAD 22.31 - - - -
MR dedicated - 60.19 41.02 91.61 65.86
[…]
TAD$+$MR Alt. 21.36 64.89 43.76 92.02 66.48
Table D: Effect of task fusion learning evaluated on Charades and Charades-STA test set.
Method policy TAD MR
mAP mAP@50  R5@50  R5@70
TAD 38.60 58.31 - -
MR dedicated - - 77.28 53.86
[…]
TAD$+$MR Sync. 39.83 60.29 80.54 57.04
[…]
Table E: Effect of task fusion learning evaluated on validation set of ANet for TAD and val_2 set of ANet-Caption for MR.
Task fusion learning on Charades and ANet. We examine the impact of task fusion learning and different sampling methods on Charades (Table D) and ANet (Table E). In terms of pre-training, both Charades and ANet benefit from this approach, except for the R5@50 and R5@70 in the MR task of Charades. Regarding the effect of co-training on these two datasets, our results align with Ego4D’s findings in Sec. 6.3. The synchronized task sampling method (referred to as “Sync.”) yields significant improvements in the co-training results for both TAD and MR tasks. Specifically, Charades shows an increase of +1.75% mAP in TAD and +3.71% R1@50 in MR, whole ANet sees an increase of +1.23% mAP in TAD and +3.26% R5@50 in MR. Additionally, we explore the effect of the alternating task sampling (referred to as “Alt.”). This sampling leads to significant enhancements in MR for both datasets (+4.70% R1@50 of Charades, +3.55% R5@50 of ANet).
Figure A: Visualization of annotations from Charades and Charades-STA. The events in green presented above the videos are natural language descriptions from MR, while the actions in blue displayed below the videos belong to predefined categories of TAD.
(a) Qualitative results on Charades and Charades-STA.
(b) Qualitative results on ANet and ANet-Caption.
Figure B: Qualitative results. The qualitative results provide evidence of the effectiveness of accurate boundary regression and moment detection using natural language descriptions from both MR and TAD. The queries in green presented above the videos are natural language descriptions from MR, while the queries in blue displayed below the videos belong to predefined categories of TAD.
Appendix C Qualitative Results
This section showcases the qualitative results obtained from videos that encompass both TAD and MR tasks, as depicted in Figure B. In our work, we convert the predefined action categories of TAD into natural language queries and treat these queries the same way as queries from MR. The qualitative results demonstrate the capability of unified moment detection using natural language descriptions from both TAD and MR.

### v2

@@ Abstract
Temporal Action Detection (TAD) focuses on detecting pre-defined actions, while Moment Retrieval (MR) aims to identify the events described by open-ended natural language within untrimmed videos. Despite that they focus on different events, we observe they have a significant connection. For instance, most descriptions in MR involve multiple actions from TAD. In this paper, we aim to investigate the potential synergy between TAD and MR. Firstly, we propose a unified architecture, termed Unified Moment Detection (UniMD), for both TAD and MR. It transforms the inputs of the two tasks, namely actions for TAD or events for MR, into a common embedding space, and utilizes two novel query-dependent decoders to generate a uniform output of classification score and temporal segments. Secondly, we explore the efficacy of two task fusion learning approaches, pre-training and co-training, in order to enhance the mutual benefits between TAD and MR. Extensive experiments demonstrate that the proposed task fusion learning scheme enables the two tasks to help each other and outperform the separately trained counterparts. Impressively, UniMD achieves state-of-the-art results on three paired datasets Ego4D, Charades-STA, and ActivityNet. Our code is available at https://github.com/yingsen1/UniMD.
@@ 1 Introduction
Temporal Action Detection (TAD) and Moment Retrieval (MR) are two similar tasks that seek to identify specific events and their corresponding temporal segments within untrimmed videos. TAD focuses on identifying temporal segments related to a single action, while MR aims to identify temporal segments that align with a natural language description. These tasks have significant connections, for instance, most natural language descriptions in MR are a combination of multiple actions from TAD. However, most existing works [59, 7, 23, 61, 53, 25] treat TAD and MR as two independent tasks, employing separate models for each. In this paper, we aim to answer a question: whether the two tasks can benefit each other by fusing them using a single model?
(a) Illustration of TAD, MR, and UniMD. UniMD performs both TAD and MR in a single model (i.e. same model parameters).
(b) Co-train results evaluated on ActivityNet under various data volumes in the training set. Dedicated models are trained separately for specific tasks, using overall videos as training data.
Figure 1: Our proposed model, UniMD, can simultaneously perform TAD and MR. When co-trained using a fraction of the training data, it can even achieve superior performance compared to dedicated models, such as 25% training data in MR and 50% in TAD.
To investigate such potential synergies between TAD and MR, we propose a new task formulation termed Moment Detection (MD) that aims to address both TAD and MR tasks simultaneously, as illustrated in Fig. 1(a). For MD, we design a task-unified architecture, Unified network for Moment Detection (UniMD), with a task fusion learning approach, to enhance the performance of both tasks:
@@
(2) Task fusion learning. To enhance synergies between TAD and MR, we explore task fusion learning to promote their mutual influence. Specifically, we examine and discuss the impact of pre-training and co-training in task fusion learning. Moreover, we introduce two co-training methods: synchronized task sampling and alternating task sampling. Synchronized task sampling prioritizes video samples encompassing both tasks, ensuring each training iteration includes both tasks. In contrast, alternating task sampling updates the network based on an alternating task at each iteration. Among them, co-training with synchronized task sampling effectively enhances the synergy and leads to distinct improvements for each task. As shown in Fig. 1(b), the proposed co-trained model can achieve better results than dedicated models, even with only a subset of training data, i.e., 25% training videos for MR and 50% training videos for TAD. This demonstrates that the mutual benefits are not merely from the increased quantity of annotations, but rather from the enhanced effectiveness of co-training.
To summarize, this paper makes three contributions: (i) we propose a unified framework, termed UniMD, to tackle TAD and MR tasks simultaneously. It links actions and events by inheriting the text encoder of CLIP to encode the query and predicts the action- or event-relevant temporal boundaries via the query-dependent heads. (ii) To the best of our knowledge, we are the first to exploit task fusion learning to investigate whether the two tasks can help each other and propose an effective co-training method to enhance their synergy, which provides valuable insights and experiences for training large language-based video models (e.g. LLaMA-vid [24]) on action/event detection. (iii) Extensive experiments demonstrate that UniMD achieves state-of-the-art performance across various benchmarks: namely, 23.25% mAP in Ego4D-MQ, 14.16% R1@30 in Ego4D-NLQ, 63.98% R1@50 in Charades-STA, and 60.29% mAP@50 in ActivityNet.
@@ 2 Related Work
Temporal Action Detection (TAD). TAD is dedicated to the identification of predefined actions and their temporal segments in untrimmed videos. It can be classified into two-stage methods [57, 66, 64] and one-stage methods [59, 26, 28]. Two-stage methods initially generate video segments as action proposals, which are subsequently categorized into actions and refined in terms of their temporal boundaries. Conversely, one-stage methods aim to localize actions within a single shot without relying on action proposals. Some TAD tasks involve a substantial number of action categories densely labeled in videos. To tackle this challenge, several approaches [7, 40, 9, 8] have been developed to strengthen the dependencies between actions and capture both short-term and long-term dependencies.
Moment Retrieval (MR). Unlike TAD, MR aims to identify specific actions using open-ended natural language descriptions. MR can be categorized into one-stage and two-stage methods. Two-stage methods [14, 60, 55, 62], also known as proposal-based methods, utilize sliding windows and proposal generation models to generate a set of candidate windows. These candidates are then ranked, and the one with the highest similarity to the given description is selected as the final result. One-stage methods [56, 58, 35, 31], also called proposal-free methods, directly compute the time pair or output confidence scores for each video snippet. Since MR searches for open-ended events rather than closed-set categories, it is crucial to fully utilize the semantic information within the queries. Therefore, several methods [56, 35, 6, 61] focus on incorporating multi-modal features and enhancing cross-modal interactions to improve boundary regression.
Vision-Language Modals. Studies on the interaction of vision and languages have been ongoing for several years. Recently, some methods [39, 16, 54] scale up training processes by leveraging large-scale datasets, resulting in the progress of robust visual-text alignment. Furthermore, many efforts [48, 50] have been made to extend these methods to video-text modals. Consequently, an increasing number of researches [67, 52, 21] focus on effective techniques for fusing vision-language representations and utilizing them to establish connections across various tasks, aiming for a universal objective.
@@ 3 Problem Definition
TAD focuses on identifying temporal segments of a pre-defined set of action categories in untrimmed videos, as shown in the first row of Fig. 1(a). Formally, given $C$ action categories, a TAD function $f_{\text{tad}}$ takes video images/features $X$ as input and generates the temporal segments of the action categories:
@@
where $s$ and $e$ denote the start time and end time of the action category $c_{\text{tad}}\in[1,C]$. Different from TAD, as shown in the second row of Fig. 1(a), MR aims to predict the temporal segments of the events described by the open-ended natural language:
@@ 4 Task-unified Architecture
where $Q=\left\{q_{i}|i\in[1,C+N]\right\}$, and $conf\in\mathbb{R}^{C+N}$ denotes the foreground confidence for the close-set action categories from TAD or the events from MR.
@@ 4.1 Encoder
where $Z=\left\{Z^{1},Z^{2},\dots,Z^{L}\right\}$ and $L$ is the number of feature pyramid levels.
@@ 4.2 Decoder
$\left\{conf^{t}\mid t\in[1,T]\right\}$ $=\Phi_{\text{cls}}\left(G,Z\right)$ (7)
$=\Phi_{\text{sigmoid}}(\Phi_{\text{scale}}(G\odot\Phi_{\text{conv}}(Z))))),$
@@
$\left\{(d_{\text{s}}^{t},d_{\text{e}}^{t})\mid t\in[1,T])\right\}$ $=\Phi_{\text{reg}}(G,Z)$ (9)
$=\Phi_{\text{relu}}(\Phi_{\text{scale}}(\Phi_{\text{q-conv}}(\Phi_{\text{conv}}(Z)))),$
@@ 6.2 Implementation Details
Network. We use the text encoder from CLIP-B [39], with frozen parameters during training. The vision encoder includes seven ConvNext blocks. Within them, two blocks serve as a stem, while the remaining five layers 2$\times$ downsample to generate the feature pyramid. The FPN incorporates three layers of BiFPN [44]. In decoder, the classification head comprises two convolutional layers and one query-dependent convolutional layer. The regression head follows the same configuration, and includes a query transformation branch with three fully connected (fc) layers. More details can be found in supplementary material.
@@
Training. The training of three paired datasets utilizes a consistent AdamW optimizer and follows the same loss weights ($\lambda_{tad}=3$, $\lambda_{mr}=1$). Supplementary material will illustrate the slight differences in other settings.
@@ 6.3 Ablation Study
(a) Effect of query dependent regression head evaluated on the Ego4D validation set under individual training. We investigate the effectiveness of queries in the regression head (reg_head) and the number of fc layers in the MLP of the regression head (#$\text{\#}FC_{reg}$).
@@
(b) Effective of task fusion learning evaluated on the Ego4D validation set. For simplicity, synchronized task sampling and alternating task sampling are abbreviated as “Sync.” and “Alt.”.
@@
(c) Effect of loss weights on validation of Ego4D, using synchronized task sampling.
@@
(d) Impact of various ratios of data volumes evaluated on validation of ANet and val_2 of ANet-Caption.
In this section, we conduct ablation experiments on the proposed query dependent head, the task fusion learning, loss weights of two tasks, and data volumes in the training set, as presented in Table 1.
Query dependent regression head. As shown in Table 1(a), the regression head with queries as input yields better performance for both tasks (21.86$\rightarrow$22.61mAP in TAD and 12.49$\rightarrow$13.99R1@30 in MR). This enhancement is attributed to the semantic information of the query, which enables more precise time intervals for specific actions. In the query transformation branch of this head, the MLP that is equipped with 3 fc layers achieves the best performance.
[…]
Loss weights. The impact of different loss weights on the outcomes is significant, as indicated by Table 1(c). For example, when using a “1-1” loss weight, the co-training performance is inferior to the dedicated model for TAD (22.55 vs. 22.61 mAP). Therefore, we carefully choose a loss weight of “3-1” to ensure that the co-train models are trained with a suitable balance of tasks.
Data volumes. The impact of data volumes on the validation of ANet is presented in Table 1(d), By controlling the number of training data for each task (e.g. 0.5-0.5 means only half of the videos from TAD and MR), we observe a positive effect on both tasks with an increase in data volumes, as shown in Fig. 1. Noteworthily, in TAD, the co-trained model using only 50% of the training videos slightly outperforms the dedicated model, while in MR, even using only 25% of the training videos significantly surpasses the dedicated model. It reveals that the mutual benefits are not merely derived from the increased quantity of annotations, but rather from the enhanced effectiveness of co-training.
@@ 6.4 Comparison with the state-of-the-art
Table 2: Comparison with the state-of-the-art on Ego4D validation set and test set.
@@
EgoVLP (VSGN)[27] EgoVLP 10.69 29.59 - - - -
InternVideo (AF)[5] VideoMAE-verb 20.69 37.24 - - - -
@@
EgoVLP (VSLNet)[27] EgoVLP - - 10.84 6.81 18.84 13.45
InternVideo (VSLNet)[5] VideoMAE-verb - - 12.32 7.43 21.73 15.07
@@
val UniMD+Sync. VideoMAE-verb 24.12 44.14 14.48 10.09 35.00 25.04
EgoVLP (VSGN)[27] EgoVLP 9.78 27.98 - - - -
InternVideo (AF)[5] VideoMAE-verb 19.31 35.58 - - - -
@@
test UniMD+Sync. VideoMAE-verb 23.25 44.80 14.16 10.06 26.95 19.16
Table 3: Comparison with state-of-the-art methods evaluated on test set of Charades for TAD and Charades-STA for MR.
@@
MomentDETR[23] SLOWFAST+CLIP - 53.63 31.37 - -
MomentDETR (pt)[23] SLOWFAST+CLIP - 55.65 34.17 - -
QD-DETR[34] SLOWFAST+CLIP - 57.31 32.55 - -
@@
Table 4: Comparison with state-of-the-art on the validation set of ANet for TAD and the val_2 set of ANet-Caption for MR. The features used in PRN cannot be obtained.
@@ 7 Conclusion
This paper provides an answer to the question of “whether the TAD and MR tasks can benefit each other by fusing them into a single model?”. To achieve this, we begin by designing a task-unified network, termed UniMD, with uniform interfaces for task input and output. We further explore various task fusion learning methods to enhance the collaboration between TAD and MR. Through experiments on three paired datasets, we present evidence that the task fusion learning approach effectively improves the performance of both tasks.